# Batch Predictive Inference

### v4

@@ 2.2 Computationally tractable examples of batch PI
We now turn to discussing how the batch PI procedure simplifies to become computationally tractable in special cases of interest.
@@ 2.2.1 Inference on a quantile
This corresponds to the batch score function $h:(s_{1},s_{2},\ldots,s_{m})\mapsto s_{\zeta}$ in Condition 1.
Observe that for this special case, we have full access to the ordering of $h$ values without knowing the exact score values, i.e., we know $S_{(i_{1})}\leqslant S_{(i_{2})}$ when $i_{1}\leqslant i_{2}$, even if the actual values of $S_{(i_{1})}$ and $S_{(i_{2})}$ are unknown. Therefore, denoting by $r_{(\zeta)}$ the $\zeta$-th smallest element in $r=(r_{1},\ldots,r_{m})$, we can set $\tilde{h}(r_{1},\cdots,r_{m})=r_{(\zeta)}$. This choice of $\tilde{h}$ recovers the exact ordering of $h$ values, in the sense that $h(S_{(r_{1})},\ldots,S_{(r_{m})})\leqslant h(S_{(r^{\prime}_{1})},\ldots,S_{(r^{\prime}_{m})})$ if and only if $\tilde{h}(r_{1},\cdots,r_{m})\leqslant\tilde{h}(r^{\prime}_{1},\cdots,r^{\prime}_{m})$. Thus, as per our discussion from Remark 1, this choice of $\tilde{h}$ is “optimal” in a sense.
Then, $q_{L}$ and $q_{U}$ in (6) can be expressed as
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\cdot\delta_{r_{(\zeta)}}\right),\qquad q_{U}=Q_{1-\gamma}\left(\sum_{r=r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\cdot\delta_{r_{(\zeta)}}\right).$
Further, since |$|\{r\in H:r_{(\zeta)}=k\}|=_{k}\mathrm{H}_{\zeta-1}\cdot_{n-k+2}\mathrm{H}_{m-\zeta}=\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}$ for $k\in[n+1]$, we have the following explicit expressions:
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right),q_{U}=Q_{1-\gamma}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right).$ (9)
Next, observe that $B_{U}$ in (7) for this setting can be simplified as
$B_{U}=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}=\max\left\{S_{r_{(\zeta)}}:r\in H,r_{(\zeta)}\leqslant q_{U}\right\}=S_{(q_{U})},$
and similarly $B_{L}=S_{(q_{L}-1)}$. Therefore, batch PI reduces to the following $(1-\alpha)$-prediction set for $S_{(\zeta)}^{\textnormal{test}}$:
@@
If the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable, the prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})$ from (9) and (10) satisfies $\mathbb{P}\left\{{S_{(\zeta)}^{\textnormal{test}}\in\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$.
For this procedure, the computational cost arises only from computing $q_{L}$ and $q_{U}$, and is relatively low since they are quantiles of discrete distributions with support size $n+1$.
@@
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k-1}\right),\quad q_{U}=Q_{1-\gamma}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right).$
Step 2: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$.
Return: Prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})=\left[S_{(q_{L})},S_{(q_{U})}\right]$
@@
In Appendix C, we extend the above idea to describe the simplification of the batch PI procedure for general sparse functions $h$, where $h(s_{1},\cdots,s_{m})$ depends only on a small number of the $s_{j}$s.
@@ 2.2.2 Inference on the mean and general compositionally structured functions
To introduce our algorithm and ideas, let us first consider the simpler case where function $\tilde{h}$ is the sum, $\tilde{h}(r_{1:m})=\sum_{j=1}^{m}r_{j}$, for all $r_{1:m}\in H$. This is equivalent to the mean, up to scale. In that case, the problem becomes to find the number—denoted $C_{m,n,k}$—of the positive integer solutions $r_{1:m}=(r_{1},\ldots,r_{m})$ to the equation $r_{1}+r_{2}+\ldots+r_{m}=k$ with $1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n$. Once we have $C_{m,n,k}$, we can simplify $q_{U}$ to
$q_{U}=Q_{1-\gamma}\left(\sum_{k\in\text{range}(\tilde{h})}\frac{C_{m,n,k}}{\binom{n+m}{m}}\delta_{k}\right).$ (11)
The key idea is to compute the numbers $C_{m,n,k}$ via dynamic programming. In particular, we find a recursion on the number of solutions depending on the value of $r_{m}\in[n]$. Consider any $a\in\{1,\ldots,n\}$. For a solution $r_{1:m}$, if $r_{m}=a$, then $r_{1}+\ldots+r_{m-1}=k-a$. By definition, there are $C_{m-1,n,k-a}$ such solutions. Thus, by considering all possible values of $a$ for $r_{m}\in[n]$, we obtain the recursion $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,n,k-a}.$
This can be used to develop a dynamic programming algorithm to find $C_{m,n,k}$, by sequentially computing all relevant values of $C_{1,*,*}$, $C_{2,*,*}$, …, $C_{m,*,*}$, where * denotes omitted indices. The initial conditions are: $C_{1,n,k}=I(1\leqslant k\leqslant n)$. Further, note that $C_{m,n,k}=0$ if $k<m$; because we need $m\leqslant r_{1}+r_{2}+\ldots+r_{m}=k$ to have a solution. Hence, the range or the recursion can be limited slightly, see Algorithm 3. We also point out that $\sum_{a=1}^{b}\cdot$ is considered to be zero when $b<1$.
  Input: number of summands $m$, maximum summand $n$, target sum $k$
  Initialize $C_{1,\tilde{n},\tilde{k}}=I(1\leqslant\tilde{k}\leqslant\tilde{n})$ for $\tilde{n}\in[n],\tilde{k}\in[k]$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{k}=1$ to $k$ do
        for $\tilde{n}=1$ to $n$ do
           $C_{\tilde{m},\tilde{n},\tilde{k}}\leftarrow\sum_{a=1}^{\min(\tilde{n},\tilde{k}-\tilde{m}+1)}C_{\tilde{m}-1,a,\tilde{k}-a}$
        end for
     end for
  end for
  Output: $C_{m,n,k}$
Algorithm 3 Computation of $C_{m,n,k}$ when the rank-ordering function $\tilde{h}$ is the sum
The running time of this algorithm is $O(mkn^{2})$ flops, due to a triple loop (each ranging from unity to $m,k,n$, respectively) and as the innermost computation takes $O(n)$ steps. Thus, since the range of $\tilde{h}$ ranges between $m$ and $(n+1)m$, computing $q_{U}$ from (11) by computing $C_{m,n,k}$ for all $k\in\text{range}(\tilde{h})$ has complexity $O(m^{2}kn^{3})$. [^3] Alternatively, for even faster computation with moderate sample sizes, one can estimate the quantiles $q_{L}$ and $q_{U}$ using sample quantiles. Specifically, drawing a sample from $H$ is equivalent to drawing $m$ samples from a uniform distribution over $[n+1]$ with replacement, allowing us to construct samples of $\tilde{h}(r_{1:m}),r_{1:m}\sim\text{Unif}(H)$. This approach leads to an accurate estimate of $q_{L}$ and $q_{U}$ if a sufficient number of samples is drawn.
More generally, suppose that for all $r\geqslant 1$, there is a strictly increasing function $\tilde{\Gamma}(\cdot;r):\{0,1,\ldots\}\to\{0,1,\ldots\}$ such that for any $\kappa\geqslant 1$,
@@
Here the function $\tilde{\Gamma}$ could be made to depend on $\kappa$, i.e., having $\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}_{\kappa}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa}),$ but we omit this for simplicity. For instance, for our previous example of the sum $\tilde{h}(s_{1:\kappa})=\sum_{j\in[\kappa]}s_{j}$, we can take $\tilde{\Gamma}(a;r)=a+r,$ for all positive integers $\kappa,r,a$. Then the same reasoning by partitioning on the possible values of $r_{m}$ yields that $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}$, where $\tilde{\Gamma}^{-1}(\cdot;a)$ denotes the inverse of the function $x\mapsto\tilde{\Gamma}(x;a)$. Here, the understanding is that if the equation $\tilde{\Gamma}(x;a)=k$ does not have a solution in $x$, then $C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}=0$.
This recursion immediately leads to a dynamic programming algorithm similar to Algorithm 3, see Algorithm 4. The initial conditions $C_{1,n,k}$ are either one or zero, depending on whether or not the equations $\tilde{\Gamma}(0;s)=k$ have a solution $1\leqslant s\leqslant n$.
  Input: Rank-ordering function $\tilde{h}$ such that for any $r\geqslant 1$, there is a strictly increasing function $\tilde{\Gamma}(\cdot;r):\{0,1,\ldots\}\to\{0,1,\ldots\}$ such that for any $\kappa\geqslant 1$, $\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa})$. Number of summands $m$, maximum summand $n$, target sum $k$
  Initialize $C_{1,\tilde{n},\tilde{k}}=I(1\leqslant\tilde{k}\leqslant\tilde{n})$ for $\tilde{n}\in[n],\tilde{k}\in[k]$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{k}=1$ to $k$ do
        for $\tilde{n}=1$ to $n$ do
        end for
     end for
  end for
  Output: $C_{m,n,k}$
Algorithm 4 Computation of $C_{m,n,k}$ for a general compositional rank-ordering function $\tilde{h}$
Computation of endpoints. The computation of the interval endpoints $B_{L},B_{U}$ from (7) can be performed efficiently in a similar way. For concreteness, we consider $B_{U}$, and the reasoning for $B_{L}$ is entirely analogous. Now, $B_{U}=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}.$
For illustration, we will again first consider the case where $h(S_{(r_{1})},\ldots,S_{(r_{m})})=S_{(r_{1})}+S_{(r_{2})}+\ldots+S_{(r_{m})}$ and $\tilde{h}(r_{1:m})=r_{1}+\ldots+r_{m}$ for all $r_{1:m}$, $S_{(r_{1})},\ldots,S_{(r_{m})}$. The problem becomes to compute
As above, we can obtain a recursion by considering the possible values of $r_{m}$, to find that $M_{m,n,q}=\max\{M_{m-1,a,q-a}\mid 1\leqslant a\leqslant\min(n,q-m+1)\}$. This recursion can be initialized with $M_{1,n,q}=S_{\min(n,q)}$, leading to a similar dynamic programming algorithm (Algorithm 5).
  Input: number of summands $m$, maximum summand $n$, upper bound on sum $q$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{q}=1$ to $q$ do
        for $\tilde{n}=1$ to $n$ do
        end for
     end for
  end for
  Output: $M_{m,n,q}$
Algorithm 5 Computation of $M_{m,n,q}$
More generally, consider a finite set $\mathcal{H}\subset\mathbb{R}$ containing, for all $\kappa\in[m]$, the values $h(S_{(r_{1})},\ldots,S_{(r_{\kappa})})$ $1\leqslant r_{1}\leqslant\ldots\leqslant r_{\kappa}\leqslant n$. Suppose that (12) holds, and that similarly, for all $r\geqslant 1$, there is a strictly increasing function $\Gamma(\cdot;r):\mathcal{H}\to\mathcal{H}$ such that for any $\kappa\geqslant 1$,
@@
If the calibration and test set sizes are very large, the above algorithms can still have a high cost. However, in certain cases of interest, especially for the central case of the mean, a straightforward procedure for inference is based on concentration inequalities. For instance, if $Y\in[a,b]$ almost surely, then by Hoeffding’s inequality, the prediction set
@@
has $(1-\alpha)$ coverage for the mean of test outcomes. Thus, very large sample sizes can be handled with concentration inequalities, while for moderate sample sizes, our algorithms remain computationally efficient, whereas the concentration-based method may result in trivial prediction sets.
2.3 Simultaneous inference on multiple quantiles
So far, we have proposed methods for univariate target functions $g$ or $h$. In this section, we extend the idea of batch PI to provide a simultaneous prediction set for multiple quantiles of the scores, e.g., $h(s_{1},\cdots,s_{m})=(s_{\zeta_{1}},s_{\zeta_{2}})^{\top}$ for all $s_{1:m}$. This will allow us to provide fine-grained control of the test distribution, for instance by obtaining a prediction set for the interquartile range.
$\begin{split}L(\rho_{1},\ldots,\rho_{l})&:=\left|\left\{(r_{1},\ldots,r_{m})\in H:r_{t_{1}}=\rho_{1},\ldots,r_{t_{l}}=\rho_{l}\right\}\right|\\ &=_{\rho_{1}}\mathrm{H}_{t_{1}-1}\cdot\left[\prod_{j=1}^{n}{}_{\rho_{j+1}-\rho_{j}+1}\mathrm{H}_{t_{j+1}-t_{j}-1}\right]\cdot_{n-\rho_{l}+2}\mathrm{H}_{m-t_{l}}.\end{split}$ (14)
$F_{n,m}(w_{1},w_{2},\cdots,w_{l};q_{1},q_{2},\cdots,q_{l})$ =|$=\left|\left\{(r_{1},r_{2},\cdots,r_{m})\in H,w_{j}\leqslant r_{t_{j}}\leqslant q_{j},\;\forall\;j\in[l]\right\}\right|$
$=\sum_{\rho_{1}=w_{1}}^{q_{1}}\sum_{\rho_{2}=\max\{\rho_{1},w_{2}\}}^{q_{2}}\cdots\sum_{\rho_{m}=\max\{\rho_{m-1},w_{l}\}}^{q_{l}}L(\rho_{1},\cdots,\rho_{l}).$ (15)
Theorem 2.
Suppose that the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable, and that $(w_{1},w_{2},\cdots,w_{l})$ and $(q_{1},q_{2},\cdots,q_{l})$ satisfy $F_{n,m}(w_{1},w_{2},\cdots,w_{l};q_{1},q_{2},\cdots,q_{l})\geqslant(1-\alpha)\cdot\binom{n+m}{m}.$ Let $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$. Then
We also mention that Gazin et al. (2024) provided an approach that they refer to as ‘template’, which could also be used to derive joint prediction sets for the order statistics of the test scores.
Continuing with our argument, it remains to determine vectors $(w_{1},\cdots,w_{l})$ and $(q_{1},\cdots,q_{l})$ that satisfy the condition of Theorem 2. For instance, we can consider the following procedure. Let $\tilde{t}_{j}=\text{round}(t_{j}\cdot n/m)$ for $j\in[l]$ represent—roughly speaking—the expected rank of the $j$-th largest test score among the $n$ calibration scores. Then our idea is to center the indices $w_{j}=\tilde{t}_{j}-a$, $q_{j}=\tilde{t}_{j}+a$, $a\geqslant 0$, around $\tilde{t}_{j}$, for $j\in[l]$. Then, we find the smallest $a\in\mathbb{N}$ such that
$F_{n,m}\left((\tilde{t}_{1}-a)\vee 1,\cdots,(\tilde{t}_{l}-a)\vee 1;(\tilde{t}_{1}+a)\wedge(n+1),\cdots,(\tilde{t}_{l}+a)\wedge(n+1)\right)\\ \geqslant(1-\alpha)\binom{n+m}{m},$
and denote it by $t$. Then define
$L=(S_{((\tilde{t}_{1}-t-1)_{+})},\cdots,S_{((\tilde{t}_{2}-t-1)_{+})}),\quad U=(S_{(\min\{\tilde{t}_{l}+t,n+1\})},S_{(\min\{\tilde{t}_{2}+t,n+1\})},\cdots,S_{(\min\{\tilde{t}_{l}+t,n+1\})}).$ (16)
Applying Theorem 2, we have the following result.
Corollary 2.
In Section 4.3, we provide experimental results for the specific case of inference on quartiles $S_{(\text{round}(0.25m))}^{\textnormal{test}},S_{(\text{round}(0.75m))}^{\textnormal{test}}$ with the following guarantee: $\mathbb{P}\left\{{L\leqslant S_{(\text{round}(0.25m))}^{\textnormal{test}}\leqslant S_{(\text{round}(0.75m))}^{\textnormal{test}}\leqslant U}\right\}\geqslant 1-\alpha$. For clarity, we include the specific procedure for this task below.
Step 2: Compute $t=\min\left\{a\in\mathbb{N}:\sum\limits_{\rho_{1}=\max\{\tilde{t}_{1}-a,1\}}^{\min\{\tilde{t}_{2}+a,n+1\}}\sum\limits_{\rho_{2}=\rho_{1}}^{\min\{\tilde{t}_{2}+a,n+1\}}{}_{\rho_{1}}\mathrm{H}_{t_{1}-1}\cdot{}_{\rho_{2}-\rho_{1}+1}\mathrm{H}_{t_{2}-t_{1}-1}\cdot_{n-\rho_{2}+2}\mathrm{H}_{m-t_{2}}\geqslant(1-\alpha)\cdot\binom{n+m}{m}\right\}$.
Return: Bounds $L=S_{(\max\{\tilde{t}_{1}-t-1,0\})}$ and $U=S_{(\min\{\tilde{t}_{2}+t,n+1\})})$.
Algorithm 6 Batch Predictive Inference for quartiles

### v5

@@ 2.3 Computationally tractable examples of batch PI
We now turn to discussing how the batch PI procedure simplifies to become computationally tractable in examples of interest.
@@ 2.3.1 Inference on a quantile
This problem has many motivations, for instance in predicting tail events. Consider for instance predicting the 95th percentile of the stock returns among several stocks. This becomes a problem of predictive inference on the quantiles. Similarly, if we are interested in the median of the hours of sunshine or rain levels over the next few days (or locations, etc), this is a problem of predictive inference on the quantiles.
Formally, inference on $S_{(\zeta)}^{\textnormal{test}}$ corresponds to the batch score function $h:(s_{1},s_{2},\ldots,s_{m})\mapsto s_{\zeta}$ in Condition 1. Observe that for this special case, we have full access to the ordering of $h$ values without knowing the exact score values, i.e., we know $S_{(i_{1})}\leqslant S_{(i_{2})}$ when $i_{1}\leqslant i_{2}$, even if the actual values of $S_{(i_{1})}$ and $S_{(i_{2})}$ are unknown. Therefore, denoting by $r_{(\zeta)}$ the $\zeta$-th smallest element in $r=(r_{1},\ldots,r_{m})$, we can set $\tilde{h}(r_{1},\ldots,r_{m})=r_{(\zeta)}$. This choice of $\tilde{h}$ recovers the exact ordering of $h$ values, i.e.,
$h(S_{(r_{1})},\ldots,S_{(r_{m})})\leqslant h(S_{(r^{\prime}_{1})},\ldots,S_{(r^{\prime}_{m})})\;\text{ if and only if }\;\tilde{h}(r_{1},\ldots,r_{m})\leqslant\tilde{h}(r^{\prime}_{1},\ldots,r^{\prime}_{m}).$
Thus, as per our discussion from Remark 1, this choice of $\tilde{h}$ is “optimal” in a sense. Then, by observing [^9] Here $p_{n,m,\zeta}$ is the probability mass function of the $\zeta$-th order statistic from a random sample of size $m$ drawn without replacement from a finite population of size $n+m$ (e.g., Wilks, 1962, p. 243).
$p_{n,m,\zeta}(k):=\frac{|\{r\in H:r_{(\zeta)}=k\}|}{|H|}=\frac{{}_{k}\mathrm{H}_{\zeta-1}\cdot_{n-k+2}\mathrm{H}_{m-\zeta}}{{}_{n+1}\mathrm{H}_{m}}=\frac{\tbinom{k+\zeta-2}{\zeta-1}\tbinom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{{m}}}$
for $k\in[n+1]$, we have the following explicit expressions:
$q_{L}=Q_{\beta}^{\prime}\bigg{(}\sum\nolimits_{k=1}^{n+1}p_{n,m,\zeta}(k)\cdot\delta_{k}\bigg{)},\quad q_{U}=Q_{1-\gamma}\bigg{(}\sum\nolimits_{k=1}^{n+1}p_{n,m,\zeta}(k)\cdot\delta_{k}\bigg{)}.$ (13)
Next, observe that $B_{U}$ in (11) for this setting can be simplified as $B_{U}=S_{(q_{U})}$, and similarly $B_{L}=S_{(q_{L}-1)}$. Therefore, batch PI reduces to the following $(1-\alpha)$-prediction set for $S_{(\zeta)}^{\textnormal{test}}$:
@@
If the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable, the prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})$ from (13) and (14) satisfies $\mathbb{P}\{S_{(\zeta)}^{\textnormal{test}}\in\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})\}\geqslant 1-\alpha.$ Furthermore, if the scores $(S_{i})_{i\in[n+m]}$ are all distinct almost surely, the following holds:
$\mathbb{P}\{S_{(\zeta)}^{\textnormal{test}}\in\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})\}\leqslant 1-\alpha+\varepsilon_{n,m,\zeta},\text{ where }\varepsilon_{n,m,\zeta}=\max_{k\in[n+1]}p_{n,m,\zeta}(k)=O(\tfrac{1}{n}).$
Above, we additionally obtain an upper bound on the coverage for inference on quantiles. This equals $1-\alpha+\frac{1}{n+1}$ when $m=\zeta=1$—i.e., the above result recovers the guarantee for the standard conformal prediction when the test size is one. For this procedure, the computational cost arises only from computing $q_{L}$ and $q_{U}$, and is relatively low, since these are quantiles of discrete distributions with support size $n+1$.
Moreover, in this case, we can also show an optimality result. Consider prediction sets of the form $\{y:s(x,y)\leqslant S_{(r)}\}$, where $r\in[n+1]$. Indeed, in the simpler case of standard conformal prediction, it is known that all prediction sets of the form $\widehat{C}_{n}(x)=\{y\in\mathcal{Y}:s(x,y)\leqslant f(S_{1},\ldots,S_{n})\}$ that have distribution-free coverage and where $f$ is a symmetric function are of this form (Robbins, 1944). Thus, the focus on such prediction sets is not restrictive. Now, based on the arguments in Section 2.2, the coverage rate of a prediction set of this form is equal to $\mathbb{P}\left\{{R_{(n+m_{\delta})}\leqslant r}\right\}$, and the batch PI method finds the smallest $r$ such that this probability is at least $1-\alpha$, based on the exact distribution of $R_{(n+m_{\delta})}$. This also leads to a tight upper bound, and implies that it dominates any other prediction set of the form $\{y:s(x,y)\leqslant S_{(r)}\}$ that achieves valid coverage.
Proposition 2 (Optimality of batch PI for the quantile).
Consider constructing prediction sets $\widehat{C}_{n}(x)=\{y\in\mathcal{Y}:s(x,y)\leqslant S_{(r)}\}$ for some fixed rank $r\in[n+1]$ for the quantile $S_{(\zeta)}^{\textnormal{test}}$ of the test datapoints, where $S_{i}=s(X_{i},Y_{i})$ are the nonconformity scores computed on the calibration data. Among all such procedures satisfying the distribution-free guarantee $\mathbb{P}\{S_{(\zeta)}^{\textnormal{test}}\in\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})\}\geqslant 1-\alpha$ under exchangeability, the batch PI procedure yields the shortest prediction sets.
@@
[…]
Step 2: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$; denote $S_{(n+1)}=+\infty$.
Return: Prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})=\left[S_{(q_{L}-1)},S_{(q_{U})}\right]$
@@
In Section D, we extend the above method to describe the simplification of the batch PI procedure for general sparse functions $h$, where $h(s_{1},\ldots,s_{m})$ depends only on a small number of the $s_{j}$s.
@@ 2.3.2 Inference on the mean and general compositionally structured functions
To introduce our algorithm and ideas, let us first consider the simpler case where the function $\tilde{h}$ is the sum, $\tilde{h}(r_{1:m})=\sum_{j=1}^{m}r_{j}$, for all $r_{1:m}\in H$. This is equivalent to the mean, up to scale. In that case, the problem becomes to find the number—denoted $C_{m,n,k}$—of the positive integer solutions $r_{1:m}=(r_{1},\ldots,r_{m})$ to the equation $r_{1}+r_{2}+\ldots+r_{m}=k$ with $1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n$. These are known as the number of partitions of $k$ with at most $m$ parts, each of size at most $n$ (Stanley, 2011), and efficient recursive algorithms are known for computing them. Once we have $C_{m,n,k}$, we can simplify $q_{U}$ to $q_{U}=Q_{1-\gamma}\Big{(}\sum_{k\in\text{range}(\tilde{h})}\frac{C_{m,n,k}}{\binom{n+m}{m}}\delta_{k}\Big{)}.$
For pedagogical purposes, we first present the idea for computing these numbers for the mean. Consider any $a\in[n]$. For a solution $r_{1:m}$, if $r_{m}=a$, then $r_{1}+\ldots+r_{m-1}=k-a$. By definition, there are $C_{m-1,n,k-a}$ such solutions. Thus, by considering all possible values of $a$ for $r_{m}\in[n]$, we obtain the recursion $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,n,k-a}.$
More generally, consider finding the number of $1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n$ such that $\tilde{h}(r_{1:m})=k$. Suppose that for all $r\geqslant 1$, there is a strictly increasing function $\tilde{\Gamma}(\cdot;r):\{0,1,\ldots\}\to\{0,1,\ldots\}$ such that for any $\kappa\geqslant 1$,
@@
Here the function $\tilde{\Gamma}$ could be made to depend on $\kappa$, i.e., having $\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}_{\kappa}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa}),$ but we omit this for simplicity. For instance, for our previous example of the sum, $\tilde{h}(s_{1:\kappa})=\sum_{j\in[\kappa]}s_{j}$, we can take $\tilde{\Gamma}(a;r)=a+r,$ for all positive integers $\kappa,r,a$. Then the same reasoning by partitioning on the possible values of $r_{m}$ yields that $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}$, where $\tilde{\Gamma}^{-1}(\cdot;a)$ denotes the inverse of the function $x\mapsto\tilde{\Gamma}(x;a)$. Here, the understanding is that if the equation $\tilde{\Gamma}(x;a)=k$ does not have a solution in $x$, then $C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}=0$.
This recursion immediately leads to a dynamic programming algorithm similar to the one for the mean. For all algorithms mentioned in this section, see Appendix F. The initial conditions $C_{1,n,k}$ are either one or zero, depending on whether or not the equations $\tilde{\Gamma}(0;s)=k$ have a solution $1\leqslant s\leqslant n$.
The running time of this algorithm is $O(mkn^{2})$ flops, due to a triple loop (each going up to $m,k,n$, respectively) and as the innermost computation takes $O(n)$ steps. Thus, since the range of $\tilde{h}$ ranges between $m$ and $(n+1)m$, computing $q_{U}$ by computing $C_{m,n,k}$ for all $k\in\text{range}(\tilde{h})$ has complexity $O(m^{2}kn^{3})$. [^10] Alternatively, for even faster computation with moderate sample sizes, one can estimate the quantiles $q_{L}$ and $q_{U}$ using sample quantiles. Specifically, drawing a sample from $H$ is equivalent to drawing $m$ samples from a uniform distribution over $[n+1]$ with replacement, allowing us to construct samples of $\tilde{h}(r_{1:m}),r_{1:m}\sim\text{Unif}(H)$. This approach leads to an accurate estimate of $q_{L}$ and $q_{U}$ if a sufficient number of samples is drawn.
The computation of the interval endpoints $B_{L},B_{U}$ from (11) can be performed efficiently in a similar way (see Appendix F).
@@
If the calibration and test set sizes are very large, the above algorithms can still have a high cost. However, in certain cases of interest, especially for the central case of the mean, a straightforward procedure for inference is based on concentration inequalities. For instance, if $Y\in[a,b]$ almost surely, then by McDiarmid’s inequality, the prediction set
@@
has $(1-\alpha)$ coverage for the mean of test outcomes, under the i.i.d. assumption. Thus, very large sample sizes $n,m$ can be handled with concentration inequalities, while for moderate sample sizes, our algorithms remain computationally efficient—under the weaker assumption of exchangeability—whereas the concentration-based method may result in trivial prediction sets. In Section 4.3, we provide experimental results comparing the performance of the batch PI-based method and the concentration-based method.